# Modelling, Design Optimization and Prototype development of Knee Exoskeleton

### v1

@@ Abstract
This study focuses on enhancing the design of an existing knee exoskeleton by addressing limitations in the range of motion (ROM) during Sit-to-Stand (STS) motions. While current knee exoskeletons emphasize toughness and rehabilitation, their closed-loop mechanisms hinder optimal ROM, which is crucial for effective rehabilitation. This research aims to optimize the exoskeleton design to achieve the necessary ROM, improving its functionality in rehabilitation. This can be achieved by utilizing kinematic modeling and formulation, the existing design was represented in the non-linear and non-convex mathematical functions. Optimization techniques, considering constraints based on human leg measurements, were applied to determine the best dimensions for the exoskeleton. This resulted in a significant increase in ROM compared to existing models. A MATLAB program was developed to compare the ROM of the optimized exoskeleton with the original design.
To validate the practicality of the optimized design, analysis was conducted using a mannequin with average human dimensions, followed by constructing a cardboard dummy model to confirm simulation results. The STS motion of an average human was captured using a camera and TRACKER software, and the motion was compared with that of the dummy model to identify any misalignments between the human and exoskeleton knee joints. Furthermore, a prototype of the knee joint exoskeleton is being developed to further investigate misalignments and improve the design. Future work includes the use of EMG sensors for more detailed analysis and better results.
@@ 1 Introduction
Exoskeleton is widely used for rehabilitation purposes. To ensure the safety the exoskeleton must align perfectly with anatomical joints to avoid potential harm, and their mechanical links and joints should work effectively within the workspace of human anatomical. To ensure this, the most common way is to find the optimal solutions of given design. In the field of exoskeleton, most design problem are complex. Depending on the objective and design of the exoskeleton, different techniques exist and some have been developed to get an approximate solution rather than global solution. Researchers have used different optimization method to mechanically design exoskeleton which fulfil their specific goal.
Zakaryan et al. [1] utilized a weight-sum-based Differential Evolution DE [2] to determine the optimal joint weights and links for their arm exoskeleton designed for rehabilitation purposes. The approach aimed to mimic the natural muscular system of human limbs by minimizing the total mass of the device, the maximal magnitude of cable tensions, and the maximal difference between magnitudes of agonist-antagonist cable tensions. Tian et al. [3] used Particle Swarm Optimization [4] to determine the ideal link lengths that reduce the leg exoskeleton’s force transfer and support the user from a sitting to a standing position. In arm support exoskeleton, Du Z. et al. [5] used Particle Swarm Intelligence technique to obtain the optimal positioning of rope within cable driven mechanism for maximum actuating torque and the optimized link length for minimum alignment between robot and human.
@@
Li et al. [6] used NSGA-II to determine optimal link lengths for a hand exoskeleton, maximizing force transmission and minimizing contact force differences and ejection. Lee et al. [7] optimized rotational angles and joint distribution for a wrist exoskeleton, enhancing accuracy and dexterity using NSGA-II. Hunt et al. [8] identified the best actuator location to maximize shoulder exoskeleton stiffness. Deboer et al. [9] used NSGA-II to find optimal link lengths, spring stiffness, angles, and displacements for a leg exoskeleton, reducing peak power and total length. Paez et al. [10] optimized link lengths and joint locations for a knee exoskeleton to minimize joint load and torque deviation, promoting natural motion.
Rituraj et al. [11] determined the optimal link lengths and angles for their assistive/rehabilitative knee exoskeleton for rehabilitation and assistive use. The optimal configuration minimizes the maximum distance between the actuators and the load on the device. According to Ttchiersky et al. [12], the researchers were able to determine the optimal placement of actuators by Genetic Algorithm (GA) for their shoulder exoskeleton, resulting in the achievement of maximum force transmission for assistive use. Du J. et al. [13] used GA to obtained the optimal link lengths to maximize the force transmission of their hand exoskeleton for rehabilitation
McDaid [14] used WBGA to find optimal link lengths for a leg exoskeleton, aiming to maximize workspace, avoid singularities, and minimize size. Asker et al. [15] employed WBGA to optimize link lengths for knee exoskeletons, focusing on maximizing force transmission and minimizing misalignment. Yu et al. [16] used WBGA to minimize displacement and dynamics in a knee exoskeleton’s hydraulic cylinder. Yoon et al. [17] applied WBGA to determine optimal joint locations for a shoulder exoskeleton, reducing misalignment and frame protrusion.
Bianchi et al. [18] used LMA to optimize link lengths for a hand exoskeleton, minimizing torque during rehabilitation. Amirpour et al. [19] applied LMA to optimize link lengths for a hand exoskeleton in haptic applications, focusing on minimizing workspace dexterity differences and force angle deviations. Malizia et al. [20] used NMSM to find optimal spring locations, lengths, and angles for a leg exoskeleton, reducing torque error. Qin et al. [21] employed GAM to optimize link lengths for finger joints, minimizing misalignments in a hand exoskeleton for rehabilitation.
Researcher are still working on the Optimization of mechanical design of exoskeleton and as per author’s knowledge and given literature review, no work has been found where a knee joint exoskeleton has been optimized using interior point technique. As per author’s best knowledge, the mechanical design optimization of Knee joint exoskeleton area has been less explored. In this paper author used Interior-Point Methods Optimization technique to mechanical model the Linear Actuator based Knee joint exoskeleton for maximum range of motion during Sit to Stand (STS) motion which is discussed further in optimization section. Thus, the present work proposes the optimization of existing knee exoskeleton that is specifically design to aid in the stand-to-sit (STS) motions. It consists of kinematic modelling of the existing four bar closed loop mechanism based exoskeleton. The mathematical formulation of design has been performed to get the required objective function. The interior-point optimization technique used to get the dimensions of the exoskeleton for maximum range of rotation. Furthermore, the geometrical simulation has been performed for optimized exoskeleton.
The paper is structured as follows: Section 2 describe the architecture of existing exoskeleton and further its kinematic modelling and optimization. Section 3 discusses the experimental results for validating the STS motion. Finally, Section 4 presents the conclusions and the future scope of work.
2 Architecture of Existing Knee joint exoskeleton
Figure 1: Model of existing exoskeleton [22]
The figure 1 shows the knee exoskeleton designed by Jain et al. [22] consist of EMG sensor-based planar four-bar mechanism that is actuated by a linear DC motor. To evaluate the efficacy of the proposed design, the exoskeleton is integrated onto an artificial limb, and modeling and simulation studies are conducted using the bond graph technique across varying loads. Subsequently, a experimental investigations are performed to validate the desired torque generated by the actuator at the knee joint for performing the STS motion. Additionally, the dynamic simulation and experimental results indicate that the actuator force varies throughout the desired activity rather than remaining constant. Consequently, the assistive torque developed by the actuator at the knee joint of the exoskeleton is found to be suitable for providing optimal support to the wearer, necessitating minimal effort on their part to perform the stand-sit-stand motions. The CAD model for the knee exoskeleton was produced utilizing SolidWorks, as illustrated in Figure 1. This design is inspired by the work of Kim et al. [23] The proposed design is based on a planar four-bar mechanism actuated by a linear DC motor. It consist of universal design that allows it to be worn by anyone, irrespective of their leg length. The exoskeleton’s fifth joint is positioned at the wearer’s knee joint, while the shank length can be adjusted to fit the wearer’s calf length.
3 Kinematic Modelling
Figure 2: Schematic Diagram of existing knee joint exoskeleton
@@ 3 Kinematic Modelling
The Figure 2 describes a robotic exoskeleton intended to support human mobility. The core component of the exoskeleton is a four-bar mechanism inspired by Jain et al. This mechanism is composed of four rigid links, namely $l_{1}$, $l_{2}$, $l_{3}$, and $l_{4}$ which are linked by four pin joints, forming a quadrilateral shape. The exoskeleton’s motion is driven by a linear DC motor, which, when activated, propels a piston that sets in motion a slewing bar ($l_{2}$), causing the thigh link ($l_{6}$) to pivot around a specific point (5) while the shank link ($l_{6}$) remains attached to the ground (ankle). The design of the exoskeleton is based on a crank-rocker mechanism that meets Grashof’s criteria. This mechanism is characterized by a sum of the shortest and longest links ($l_{3}$ and $l_{4}$) that is less than the sum of the other two links ($l_{1}$ and $l_{2}$). Additionally, the link adjacent to the shortest link ($l_{4}$) remains fixed during STS motion. As a result of these features, the shortest link ($l_{3}$) acts as a crank, driving the mechanism’s motion. Meanwhile, the link next to the fixed link ($l_{1}$) behaves as a rocker, oscillating back and forth. The coordinated interplay between the links and the motor enables the exoskeleton to move its limbs, thus emulating the natural gait of humans.
The mathematical formulation based on the above exoskeleton are:
$\theta=180^{\circ}-\beta_{1}-\beta_{2}$ (1)
$\beta_{1}=\cos^{-1}\left(\left(l_{8}^{2}+l_{3}^{2}-l_{2}^{2}\right)/2l_{8}l_{3}\right)$ (2)
$\beta_{2}=\cos^{-1}\left(\left(l_{8}^{2}+l_{4}^{2}-l_{1}^{2}\right)/2l_{8}l_{4}\right)$ (3)
@@
where $\alpha_{1}$ is the angle between link $l_{5}$ and $l_{6}$. $\alpha_{2}$ is the angle between link $l_{7}$ and $l_{2}$. $\beta_{1}$ is the angle between link $l_{3}$ and $l_{8}$. $\beta_{2}$ is the angle between link $l_{8}$ and $l_{4}$. $\theta$ is the knee joint angle. d is the stroke length of linear actuator.
@@ 4.1 Problem Formulation
The angle of the knee joint in an exoskeleton is governed by a linear actuator. The maximum angle of rotation achievable by the knee joint is contingent on the minimum stroke length possible with the linear actuator. In the present setup, this minimum stroke length is 242 mm. The optimization of the six remaining links that indirectly impact the knee joint’s range of motion is essential to increase the latter. Our objective, therefore, is to maximize the objective function below to achieve this goal.
$\theta=180^{\circ}-\beta_{1}-\beta_{2}$
@@ 4.2 Design Variables
This optimization problem consists of six variables as shown below. The width of all the links are same.
$X={l_{1},l_{2},l_{3},l_{4},l_{5},l_{6}}$
Because of these six variables, the objection function must be in the form of
$f=\theta(l_{1},l_{2},l_{3},l_{4},l_{5},l_{6},d)$
@@
4.3 Constrains
In this optimization, there are fifteen constrains:
@@ 4.3 Constrains
    The four bar linkage consist of l1, l2, l3, l4 link. The four bar mechanism should follow the Grashof’s law i.e. the sum of shortest and largest link should be less than the sum of other two links:
    $l_{3}+l_{4}-l_{2}-l_{1}<0$
@@
    The above constrains follow the Grashof’s law but while designing we don’t have the information which is the shortest link in all four links:
    $l_{3}-l_{2}<0,l_{3}-l_{1}<0$
@@
    Similarly, we don’t have the information of longest link also. Therefore, we are adding below constrains:
    $l_{3}-l_{4}<0,l_{2}-l_{4}<0,l_{1}-l_{4}<0,$
@@
    The width of the thigh of human beings is always less than the length of the thigh. Therefore, we are adding below constrains:
    $l_{1}-l_{4}<0,l_{5}-l_{6}<0,l_{4}-l_{6}<0$
    Figure 3: Sit position
@@
    Since, the optimization is purely based on geometrical approach, it is important to give constrains to knee joint angle. In real design there is limit to the range of motion. To do so, we have provided singularity constrain.. The figure 3 shows the position of links in the triangle when reaching at the maximum knee joint angle of exoskeleton. The further increment of angle reduces the angle between l7 and d. When the angle between link d and l2 becomes 180 deg. then it loses its 1 degree of freedom which convert the 1 degree of freedom mechanism into structure. Therefore, the mechanism become redundant if it satisfies the constrains $l_{7}$=$l_{2}$ + $d_{min}$.
@@
    The optimized exoskeleton’s dimension can be scale up or down. To avoid these, we have set the boundaries of each variables. These boundaries are on the basis of measurement of leg dimension of average human being weighted 80 kg and 180 cm height:
    $100mm>d_{1}>50mm$
    $100mm>d_{2}>50mm$
    $100mm>d_{3}>50mm$
    $100mm>d_{4}>50mm$
    $120mm>d_{5}>50mm$
    $300mm>d_{6}>200mm$
@@ 5 Optimization method and Results
The proposed exoskeleton comprises of 6 links, and to obtain the optimal measurement for maximum range of motion for every user, synthesis of exoskeleton has been performed. The optimization of knee joint exoskeleton is a vital aspect of knee joint support. To achieve optimal outcomes, the Interior Point Method is utilized. This iterative algorithm is employed to solve convex optimization problems, unlike traditional simplex methods that traverse the edges of the feasible region. Interior point methods move through the interior of the feasible region towards the optimal solution, enabling efficient handling of large-scale problems with numerous constraints and variables. Barrier functions are utilized by interior point methods to enforce the feasibility of solutions. These functions penalize infeasible solutions, guiding the optimization process towards feasible regions.
Interior point techniques provide the ability to change the barrier according to the measurement of human limbs, and this technique presents many advantages such as efficiency, scalability, and robustness. Interior point methods exhibit polynomial-time complexity, making them suitable for large-scale optimization problems. These methods are robust against ill-conditioned problems and numerical instabilities, providing reliable solutions even in challenging scenarios. Interior point techniques scale well with problem size, enabling the optimization of high-dimensional problems encountered in modern applications such as machine learning and data analysis.
The optimization has been done using the Interior point method. This problem is non-linear inequality constrained. The interior point method which is type of barrier method doesn’t depend on the type of problem. i.e. it can be used for both for convex or non-convex problem.
@@
Design Variable Initial value (in mm) Optimum value (in mm)
@@
$l_{5}$ 80 118.63
$l_{6}$ 235 287.31
Figure 4: Optimized exoskeleton model
In order to optimize a given function, it is crucial to identify the optimal point and the corresponding minimum value. To achieve this, we adopt a conventional approach that involves stepping away from the optimal point and observing the changes that occur in the optimal solution. By doing so, we are able to obtain a more comprehensive understanding of the behaviour of the function and the impact of the surrounding environment on its performance. As an illustration, a figure 5 showcasing the function’s local minima demonstrates that the optimal values increase as we move away from the optimal point, which is situated at the centre of the x-axis. In the figure 6, the dependency of stroke length on the knee joint angle has been shown. This graph shows the non-linearly relationship between the stroke length and knee joint angle.
Figure 5: Existence of Local minima at optimal solution.
Figure 6: Knee joint angle vs Stroke length

### v3

# Optimal Kinematic Synthesis and Prototype Development of Knee Exoskeleton

@@ Abstract
The range of rotation (RoR) in a knee exoskeleton is a critical factor in rehabilitation, as it directly influences joint mobility, muscle activation, and recovery outcomes. A well-designed RoR ensures that patients achieve near-natural knee kinematics, which is essential for restoring gait patterns and preventing compensatory movements. This paper presents optimal design of one degree of freedom knee exoskeleton. In kinematic analysis, the existing design being represented by nonlinear and nonconvex mathematical functions. To obtain feasible and optimum measurement of the links of knee exoskeleton, an optimization problem is formulated based on the kinematic analysis and average human’s leg measurement. The optimized solution increases the range of motion of knee exoskeleton during sit to stand motion by $24\%$ as compared with inspired design. Furthermore, misalignment study is conducted by comparing the trajectory of human’s knee and exoskeleton’s knee during sit to stand motion. The joint movement is calculated using marker and camera system. Finally, a prototype of the knee joint exoskeleton is being developed based on optimal dimensions which validate the maximum range of motion achieved during simulation.
@@ 1 Introduction
Patients with mobility impairments resulting from stroke, osteoarthritis, or spinal cord injuries require rehabilitation for post-recovery. Knee exoskeleton have emerged as transformative tools in rehabilitation by providing mechanical support and guided movement, these devices help restore natural gait patterns, reduce muscle atrophy, and accelerate recovery [1]. However, their effectiveness depends on the refined mechanical design particularly in replicating the biomechanics of the human knee.
Among various designs, single degree-of-freedom (DoF) exoskeletons offer significant advantages over their multi-DoF counterparts, especially in terms of simplicity, reliability, and wearability. Unlike complex multi-joint exoskeletons that require advanced control systems, higher energy consumption and are susceptible to misalignment, single-DoF designs focus solely on flexion and extension, which are the primary motions needed for walking, stair climbing, and transitioning from sitting to standing. This simplified approach reduces mechanical complexity, minimizes weight, and enhances user comfort, making single-DoF exoskeletons more practical for daily rehabilitation use [2].
However, a critical challenge in single-DoF knee exoskeletons is ensuring that their joint mechanics accurately replicate the natural kinematics of the human knee. This is where a four-bar linkage mechanisms is more useful. Unlike traditional two-bar hinge structures that have a fixed Instantaneous Center of Rotation (ICR), the four-bar mechanism allows the ICR to move in a J-shaped trajectory, similar to the human knee joint [3]. By incorporating a movable pivot system—comprising femoral, tibial, and two connecting links—the four-bar design dynamically adjusts its rotation center to closely match the knee’s natural ICR trajectory.
To ensure this, the most common way is to find the optimal solutions of given design. In the field of exoskeleton, most design problem are complex. Depending on the objective and design of the exoskeleton, different techniques have been developed to get an approximate solution. Researchers have used different optimization methods to mechanically design exoskeleton which fulfill their specific goal. Zakaryan et al.[4] applied a weighted-sum Differential Evolution (DE) approach [5] to optimize joint weights and link dimensions for an arm exoskeleton, aiming to emulate human muscle behavior by minimizing mass, peak cable tensions, and imbalances between agonist and antagonist tensions. Similarly, Tian et al.[6] used Particle Swarm Optimization (PSO) to optimize link lengths for enhanced force transmission and sit-to-stand support in a leg exoskeleton. Du et al.[7] also employed PSO to fine-tune rope routing and link lengths in a cable-driven arm mechanism, focusing on maximizing torque and reducing misalignment.
@@
Deboer et al. [8] used NSGA-II to find optimal link lengths, spring stiffness, angles, and displacements for a leg exoskeleton, reducing peak power and total length. Paez et al. [9] optimized link lengths and joint locations for a knee exoskeleton to minimize joint load and torque deviation, promoting natural motion. Rituraj et al. [10] determined the optimal link lengths and angles for their assistive/rehabilitative knee exoskeleton for rehabilitation and assistive use. The optimal configuration minimizes the maximum distance between the actuators and the load on the device. McDaid [11] used WBGA to find optimal link lengths for a leg exoskeleton, aiming to maximize workspace, avoid singularities, and minimize size. Asker et al. [12] employed WBGA to optimize link lengths for knee exoskeletons, focusing on maximizing force transmission and minimizing misalignment.
Researcher are still working on the optimization of mechanical design of exoskeleton and as per author’s knowledge and given literature review, no work has been found where a knee joint exoskeleton has been optimized using interior point technique. In this paper, authors have developed an optimization algorithm based on Interior Point Method optimization technique and performed feasibility check using graphical simulation to mechanical model the Linear Actuator based Knee joint exoskeleton for maximum range of motion during Sit to Stand (STS) motion. Figure 1(b) shows the knee exoskeleton inspired by Jain et al. [13] and Kim et al. [14]. The proposed design is based on a planar four-bar mechanism actuated by a linear DC motor. It consist of universal design that allows it to be worn by anyone, irrespective of their leg length. The exoskeleton’s fifth joint is positioned at the wearer’s knee joint, while the shank length can be adjusted to fit the wearer’s calf length. The paper is structured as follows: Section 2 describe the architecture of existing exoskeleton and its kinematic modelling. Section 3 and 4 discusses the formulation of objective functions and optimization results for the STS motion. Section 5 and 6 contains the graphical simulation of optimized exoskeleton and gait analysis. Finally, Section 7 and 8 presents the validation of the optimal result by prototype development and conclusion.
2 Kinematic Modeling
(a) Schematic diagram of existing knee joint exoskeleton
(b) CAD model of knee exoskeleton.
Figure 1: Linear Actuator based Knee Exoskeleton (LAKE)
@@ 2 Kinematic Modeling
Figure 1(a) describes a robotic exoskeleton intended to support human mobility. The core component of the exoskeleton is a four-bar mechanism. This mechanism is composed of four rigid links, namely $l_{1}$, $l_{2}$, $l_{3}$, and $l_{4}$ which are linked by four pin joints, forming a quadrilateral shape. The exoskeleton’s motion is driven by a linear DC motor, which, when activated, propels a piston that sets in motion a slewing bar ($l_{2}$), causing the thigh link ($l_{6}$) to pivot around a specific point (5) while the shank link ($l_{6}$) remains attached to the ground (ankle). The design of the exoskeleton is based on a crank-rocker mechanism that meets Grashof’s criteria. This mechanism is characterized by a sum of the shortest and longest links ($l_{3}$ and $l_{4}$) that is less than the sum of the other two links ($l_{1}$ and $l_{2}$). Additionally, the link adjacent to the shortest link ($l_{4}$) remains fixed during STS motion. As a result of these features, the shortest link ($l_{3}$) acts as a crank, driving the mechanism’s motion. Meanwhile, the link next to the fixed link ($l_{1}$) behaves as a rocker, oscillating back and forth. The coordinated interplay between the links and the motor enables the exoskeleton to move its limbs, thus emulating the natural gait of humans. The mathematical formulation based on the exoskeleton in Fig 1(b) is detailed here keeping $\theta$ as the knee joint angle - representing the degree of freedom of the system.
$\theta(t)=180^{\circ}-\beta_{1}(t)-\beta_{2}(t)$ (1)
$\beta_{1}(t)=\cos^{-1}\left(\left(l_{8}^{2}+l_{3}^{2}-l_{2}^{2}\right)/2l_{8}l_{3}\right)$ (2)
$\beta_{2}(t)=\cos^{-1}\left(\left(l_{8}^{2}+l_{4}^{2}-l_{1}^{2}\right)/2l_{8}l_{4}\right)$ (3)
@@
where $d$ is the stroke length of linear actuator. $\alpha_{1}$ is the angle between link $l_{5}$ and $l_{6}$. $\alpha_{2}$ is the angle between link $l_{7}$ and $l_{2}$. $\beta_{1}$ is the angle between link $l_{3}$ and $l_{8}$. $\beta_{2}$ is the angle between link $l_{8}$ and $l_{4}$ and both $\beta$ angle changes according to design link variables explained in Section 3.2.
@@ 3.1 Problem Formulation
The angle of the knee joint in an exoskeleton is governed by a linear actuator. The maximum angle of rotation achievable by the knee joint is contingent on the minimum stroke length possible with the linear actuator. Based on Jain et al [13], this minimum stroke length is taken as 242 mm. The optimization of the six remaining links that indirectly impact the knee joint’s range of motion is essential to increase the latter. The main objective is to maximize the function $F(x)$ defined in equation (7) with design variable represented in equation (8).
$\ F(x):\theta=180^{\circ}-\beta_{1}(t)-\beta_{2}(t)$ (7)
@@ 3.2 Design Variables
The exoskeleton consists of 6 independent links whose length dimensions are considered as design variables in this problem. This selection provides a larger domain for acquiring optimal solution. Also the width of all the links are assumed to be same.
$\textbf{X}=\{l_{1},l_{2},l_{3},l_{4},l_{5},l_{6}\}^{t}$ (8)
Based on these six variables, the objection function can be represented as
$max\;f=\theta(\textbf{X},d)$
@@
3.3 Constraints
The proposed exoskeleton comprises of 6 links, and to obtain the optimal measurement for maximum range of motion for every user, constraints play a crucial role in the optimization. For optimal solution of objective function in Section 3.2, the constraints are proposed based on mechanical and human knee restrictions in real world scenario.
@@ 3.3 Constraints
    The four bar linkage consist of $l_{1}$, $l_{2}$, $l_{3}$$l_{4}$ link. The four bar mechanism should follow the Grashof’s law i.e. the sum of shortest and largest link should be less than the sum of other two links. i.e.  $l_{3}+l_{4}-l_{2}-l_{1}<0$
@@
    The above constrains follow the Grashof’s law but while designing there is no information which is the shortest link in all four links. Therefore, $l_{3}-l_{2}<0$ and $l_{3}-l_{1}<0$.
@@
    Similarly, there is no information of longest link also. Therefore, there are addition of given constrains  $l_{3}-l_{4}<0$, $l_{2}-l_{4}<0$, $l_{1}-l_{4}<0$
@@
    Since the width of the human thigh is always less than its length, a constraint is imposed as $l_{1}-l_{4}<0$, $l_{5}-l_{6}<0$, $l_{4}-l_{6}<0$ accurately reflect this anatomical relationship in the optimization problem.
@@
    Since, the optimization is purely based on geometrical approach, it is important to give constrains to knee joint angle up to the feasible limit.To achieve this, a singularity constraint has been added into the problem. Figure 2 shows the position of links in the triangle when reaching at the maximum knee joint angle of exoskeleton. The further increment of angle reduces the angle between $l_{7}$ and $d$. When the angle between link $d$ and $l_{2}$ becomes $180^{\circ}$, then it loses its one degree of freedom which convert this mechanism into structure. Therefore, the mechanism become structure if it satisfies the constrains $l_{7}$ = $l_{2}$ + $d_{min}$.
    Figure 2: Knee Exoskeleton diagram during Sit position
@@
    The optimized dimensions of the exoskeleton can be scaled up or down. To avoid these, variable founds are considered on the basis of leg dimensions of an average human having weight = 80 kg and height = 180 cm :
    $100mm>d_{1}>50mm$ , $100\;mm>d_{2}>50\;mm$ , $100\;mm>d_{3}>50\;mm$, $100\;mm>d_{4}>50\;mm$, $120\;mm>d_{5}>50\;mm$, $300\;mm>d_{6}>200\;mm$ .
@@ 4 Optimization Method and Results
The proposed exoskeleton comprises of 6 links, and to obtain the optimal measurement for maximum range of motion for every user, synthesis of exoskeleton has been performed. Interior point method is considered to solve this nonlinear inequality constrained optimization problem. The interior point method which is type of barrier method doesn’t depend on the type of problem. i.e. it can be used for both for convex or non-convex problem. The optimal solution based on this method is given in table 1. Based on the optimum result, the CAD model is prepared with mannequin to check the feasibility of knee exoskeleton and its maximum range of rotation as shown in Fig 3.
@@
Design Variable Initial Value (mm) Optimum Value (mm)
@@
$l_{5}$ 80 118.630
$l_{6}$ 235 287.310
Figure 3: (a) CAD model of optimized exoskeleton (b) Mannequin wearing exoskeleton on hospital bed
To optimize a given function, it is essential to identify both the optimal point and its corresponding minimum value. This is typically achieved by perturbing the optimal point and analyzing the resulting changes in the solution. As an illustration, a Fig 4(a) presents the function’s local minima, showing that the objective values increase as one moves away from the optimal point, which is located at the center of the x-axis. In the Fig 4(b), the dependency of stroke length on the knee joint angle has been shown. This graph shows the non-linear relationship between the stroke length and knee joint angle.
(a) Local minima at optimal solution.
(b) Knee joint angle vs Stroke length.
Figure 4: (a) Existence of local minima at optimal solution. (b) Variation of knee joint angle with respect to linear actuator’s stroke length.